# Practical Secure Delegated Linear Algebra with Trapdoored Matrices

### v2

@@ 5 A LPN Hardness Assumption and Performance Scaling
$\exists\epsilon<1$ such that fixing any $\delta>0$, there exists $n^{*}\in\mathbb{N},c>0$ such that for any $n>n^{*},n_{1}=\delta n$, $\mu=n^{\epsilon-1}$, no algorithm running in time $T=2^{O_{\delta}(n^{c})}$ can solve decisional-LPN with parameters $n,n_{1},\mu$ with advantage $\geq 2^{-\Omega_{\delta}(n^{c})}$.
@@
Both $\delta$ and $\epsilon$ directly correspond to the parameters by the same names in Theorem 3.1. In particular, slightly stronger assumptions directly translate into significant improvements to the efficiency of the algorithm. For instance, if $\delta$ is allowed to be $o(1)$, we achieve sub-multiplicative overhead for $\frac{m}{n}\geq\delta$, rather than just $m=\Omega(n)$. All terms of the form $n^{\epsilon}$ correspond to $n\mu$; if we assume that $\mu(n)=n^{\epsilon-1}$ is secure for any $\epsilon>0$, we obtain the corresponding substantial efficiency improvements. Likewise, the security of our protocol depends directly on the computational hardness assumption: a super-polynomial assumption would yield a protocol with super-polynomial security.
Note: our matrix-multiplication results are nontrivial when $\epsilon<\omega-2$. The above assumption is commonly believed to hold (nonuniformly) for all $\epsilon$, or at least for $\epsilon$ quite small.
6 A Simple Secure Delegated Matrix-Matrices Algorithm
@@
We first use LPN to construct schemes for Trapdoored-Matrix and Targeted-Trapdoored-Matrix, and therefore Secure Delegated Matrix-Matrices. These, and an associated algorithm for Secure Delegated Matrix-Matrices, are detailed in Algorithm 1. Our approach is simple: we observe that a matrix of the form $HL+S$, where $H$ and $L$ have few columns and rows respectively and $S$ is sparse, is pseudorandom by the LPN assumption and is easy to multiply by other matrices. Moreover, given a matrix $A$, once $AL^{\top}$ has been computed, it is very efficient to compute $A(LH+S)$ as $(AL)H+AS$, even if $LH+S$ represents a thin matrix (or vector).
@@
A brief note on notation: in order to aid the reader in tracking the values represented by intermediate variables, we use angled brackets around an expression, such as $\langle\mathrm{Expr}(x,y,z,\dots)\rangle$ to denote a variable whose value is equal to that of the expression $\mathrm{Expr}(x,y,z,\dots)$.
Algorithm 1 Basic Secure Delegated Matrix-Matrices via Trapdoored-Matrix and Targeted-Trapdoored-Matrix Schemes
Procedures are written for the client. The $\mathbf{Send}$ and $\mathbf{SMult}$ commands are used, respectively, to send information and to request matrix multiplications (sending their arguments if the server does not already have them) to/from the server. For efficiency, $L_{i}$s are shared over all routines.
1:procedure TM-Init
2:     Input: $m,n\in\mathbb{N},n_{1}\in\mathbb{Z}_{>0},\mu\in\mathbb{R}_{>0}$
3:     $L\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{1}\times n}$, $H\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{m\times n_{1}}$, $S\overset{{\scriptscriptstyle\$}}{\leftarrow}\mathcal{D}_{\mu}^{m\times n}$
4:     $A^{\prime}=HL+S$
5:     return $A^{\prime}$
6:procedure TM-Mult
7:     Input: $M\in R^{n\times l}$
8:     return $(SM)+(H(LM))$triangleright Equal to $A^{\prime}M$
9:procedure TTM-Init
10:     Input: $n\in\mathbb{N},n_{1}\in\mathbb{Z}_{>0},\mu\in\mathbb{R}_{>0},N\in R^{m\times n}$
11:     $L_{R}\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{1}\times n}$
12:     $P=NL_{R}^{\top}$
13:procedure TTM-Gen-Mult
14:     Input: $l\in\mathbb{N}$
15:     $H_{R}\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{l\times n_{1}}$, $S_{R}\overset{{\scriptscriptstyle\$}}{\leftarrow}\mathcal{D}_{\mu}^{l\times n}$
16:     return $(L_{R}^{T}H_{R}^{T}+S_{R}^{T},PH_{R}^{T}+NS_{R}^{T})$$\triangleright$ Equal to $B^{\prime}_{i}$ and $NB^{\prime}_{i}$, respectively
17:procedure HFMM-Init
18:     Input: $A\in R^{m\times n}$, $n_{1}\in\mathbb{Z}_{>0}$, $\mu\in\mathbb{R}_{>0}$
19:     $A^{\prime}=\textsc{TM-Init}(m,n,n_{1},\mu)$
[…]
21:procedure HFMM-Mult
22:     Input: $B_{i}\in R^{m\times l}$
23:     $\langle A^{\prime}B_{i}\rangle=\textsc{TM-Mult}(B_{i})$
24:     $(B^{\prime}_{i},\langle(A+A^{\prime})B^{\prime}_{i}\rangle)=\textsc{TTM-Gen-Mult}(l)$
25:     return $\mathbf{SMult}(\langle A+A^{\prime}\rangle,\langle B_{i}+B^{\prime}_{i}\rangle)-\langle A^{\prime}B_{i}\rangle-\langle(A+A^{\prime})B^{\prime}_{i}\rangle$
@@ 6.1 Security and Performance
The security of this scheme is a direct consequence of the LPN assumption: the matrices $A^{\prime}$ and $B^{\prime}$ that TM-Mult and TTM-Mult generate have rows or columns constructed as LPN pseudorandom vectors, and therefore $A+A^{\prime},B+B^{\prime}$ (the only information that the server receives) leak no information about $A$ and $B$.
[…]
@@
For most parameter regimes (e.g. $\delta=\Theta(1),\mu=o(1)$), the dominant cost comes from the dense multiplication $HL$, and therefore varies polynomially in subspace dimension $n_{1}$. This can be greatly reduced via LPN variants over fast matrices (e.g. sparse-LPN, Toeplitz-LPN, Quasi-cyclic-LPN [AMBD+18]), but parameter tradeoffs for these problems over large rings are not currently well-understood, and may entail worse $\mu$. In the next subsection, we present an alternate option that obtains equal or better performance (depending on the comparative sparsity assumptions of LPN and sparse-LPN) than a sparse-LPN version, but depends only on standard LPN assumptions.
In the case where $A$ is public, we can also achieve substantial improvements via dual-LPN. This simple substitution (i.e. calculating $v^{\prime}=Ps$, $Av^{\prime}=(AP)s$ for a public random square matrix $P$) takes the cost down to $O(mn\mu)$. The dual formulation will not, however, be sufficient to achieve very fast fully private matrix-matrices.
7 An Improved Secure Delegated Matrix-Matrices Algorithm
@@
The efficiency of the previous section’s algorithm is primarily determined by the $\delta$ of the LPN assumption, which is rather high. This is not (just) an artifact of our weak assumption: the security of LPN does not degrade gracefully as its parameters are reduced. In particular, once $n_{1}\mu\leq 1$, straightforward attacks from subsampling and Gaussian elimination are feasible. This increases the overhead of Algorithm 2 for large inputs. In this section, we present a more complicated algorithm that bypasses this problem to improve performance.
@@
The client cost of our earlier scheme is dominated by the costs of the multiplications $AL$ and $(AL)r_{i}$. There is a natural solution: recursively call the remote server to compute these while not leaking information about $A$ or $r_{i}$. Additionally, computations where both matrices are public (in particular, those of the form $L(B+B^{\prime})$ and similar) may be offloaded to the server. These changes result in a fairly complicated recursive protocol with several disadvantages, including a number of rounds growing exponentially in the recursion depth and substantial duplication of work, but this protocol in turn may be greatly optimized and simplified. We detail the resulting algorithm for secure delegated matrix-matrices in Algorithm 2, and spend the rest of the section demonstrating that it satisfies Theorem 3.1.
Algorithm 2 Client Routines for LPN-based Matrix-Vectors
1:procedure Mat-Mul-Check
2:     Input: $A\in R^{a\times b},B\in R^{b\times c},C\in R^{a\times c}$, $\lambda^{\prime}\in\mathbb{N}$ $W\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{c\times\lambda^{\prime}}$
3:     if $A(BW)\neq CW$ then Terminate
4:procedure Get-Checked-Cum-Prods
5:     Input: $M_{i}\in R^{a_{i}\times a_{i+1}}\forall i\in[k]$
6:     for $i\in\{2,3,\dots,k\}$ do
7:         $\langle M_{1}M_{2}\dots M_{i}\rangle=\mathbf{SMult}(\langle M_{1}M_{2}\dots M_{i-1}\rangle,M_{i})$
[…]
9:procedure Right-Hidden-Init
10:     Input: $M\in R^{s\times n}$, $n_{i}\forall i\in[d],\mu_{i}\forall i\in[d]$
[…]
12:procedure Right-Hidden-Mult
13:     Input: $V\in R^{n\times o}$.
14:     $H\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{d}\times o}$
15:     $\forall i\in[d],S_{i}\overset{{\scriptscriptstyle\$}}{\leftarrow}\mathcal{D}_{\mu_{i}}^{n_{i-1}\times o}$
16:     $V^{\prime}\leftarrow\langle L_{1}L_{2}\dots L_{d}\rangle H+\sum_{i=1}^{d}\langle L_{1}L_{2}\dots L_{i-1}\rangle S_{i}$
17:     $\langle MV^{\prime}\rangle\leftarrow\langle ML_{1}L_{2}\dots L_{d}\rangle H+\sum_{i=1}^{d}\langle ML_{1}L_{2}\dots L_{i-1}\rangle S_{i}$
[…]
19:     return $\langle M(V+V^{\prime})\rangle-\langle MV^{\prime}\rangle$.
20:procedure Hidden-Init
21:     Input: $A\in R^{m\times n}$, $n_{i}\forall i\in[d],\mu_{i}\forall i\in[d-1]$
22:     for $i\in[d]$ do $L_{i}\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{i}\times n_{i-1}}$, $S_{i}\leftarrow\mathcal{D}_{\mu_{i}}^{m\times n_{i-1}}$      $H\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{m\times n_{d}}$.
23:     for $i\in[d]$ do $\mathbf{Send}(L_{i})$
[…]
25:     $C\leftarrow[\langle L_{1}\rangle,\langle L_{1}L_{2}\rangle,\dots,\langle L_{1}L_{2}\dots L_{d}\rangle]$triangleright as a block matrix
[…]
27:     $[\langle AL_{1}\rangle,\langle AL_{1}L_{2}\rangle,\dots,\langle AL_{1}\dots L_{d}\rangle]\leftarrow\textsc{Right-Hidden-Mult}(C^{\top},A^{\top})^{\top}$triangleright Divide $C^{\top}$ into a row-vector of blocks for better efficiency
28:     $A^{\prime}=H\langle L_{1}L_{2}\dots L_{d}\rangle^{\top}+\sum_{i=0}^{d-1}S_{i+1}\langle L_{1}L_{2}\dots L_{i}\rangle^{\top}$
29:     $\mathbf{Send}(A+A^{\prime})$
30:procedure Hidden-Mult
31:     Input: $B\in R^{m\times l}$.
32:     $R\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{d}\times l}$
33:     for $i\in[d]$ do $S_{i}\leftarrow\mathcal{D}_{\mu_{i}}^{n_{i-1}\times l}$
34:     $B^{\prime}\leftarrow\langle L_{1}L_{2}\dots L_{d}\rangle R+\sum_{i=1}^{d}\langle L_{1}L_{2}\dots L_{i-1}\rangle S_{i}$
35:     $\langle AB^{\prime}\rangle\leftarrow\langle AL_{1}L_{2}\dots L_{d}\rangle R+\sum_{i=1}^{d}\langle AL_{1}L_{2}\dots L_{i-1}\rangle S_{i}$
36:     for $i\in[d]$ do $\mathbf{SMult}((L_{1}L_{2}\dots L_{i})^{\top},(B+B^{\prime}))$.
37:     $\langle A^{\prime}(B+B^{\prime})\rangle\leftarrow H\langle(L_{1}L_{2}\dots L_{d})^{\top}(B+B^{\prime})\rangle+\sum_{i=0}^{d-1}S_{i+1}\langle(L_{1}L_{2}\dots L_{i})^{\top}(B+B^{\prime})\rangle$
38:     return $\mathbf{SMult}(\langle A+A^{\prime}\rangle,\langle B+B^{\prime}\rangle)-\langle AV^{\prime}\rangle-\langle A^{\prime}(B+B^{\prime})\rangle$
7.1 Algorithm Structure
Algorithm 2 is divided into several subroutines:
  • Mat-Mul-Check: Given two matrices $A$, $B$, and a third matrix $C$, check that $AB=C$ with negligible statistical soundness error.
  • Get-Checked-Cum-Prods: Given a sequence of matrices $M_{i}\in R^{a_{i}\times a_{i+1}}$ for a nonincreasing sequence of positive integers $a_{i}$, query the server for the sequence of cumulative products $M_{1}M_{2},M_{1}M_{2}M_{3},\dots$, and check the answer up to negligible statistical soundness error.
  • Right-Hidden-Init: Initialize a simplified version of the protocol that can offload a matrix-matrix multiplication to the server while hiding one operand.
  • Right-Hidden-Mult: Apply the above protocol.
  • Hidden-Init: Initialize the full protocol.
  • Hidden-Mult: Run the initialized protocol on a target operand.
All pseudorandom matrices are defined via recursion on LPN; for instance, we observe that
$V^{\prime}$ $=L_{1}L_{2}\dots L_{d}H+\sum_{i=1}^{d}L_{1}L_{2}\dots L_{i-1}S_{i}$
$=(S_{1}+L_{1}(S_{2}+L_{2}(S_{3}+L_{3}(\dots(S_{d}+L_{d}H)))))$
That this is pseudorandom will follow from a hybrid argument on the LPN assumption: that a parenthetical is pseudorandom, combined with LPN hardness, shows that the parenthetical containing it is pseudorandom.
Recursing on LPN makes the required dense matrix-vector multiplication substantially smaller (at the cost of an additional sparse multiplication), and so improves efficiency. This mostly decouples instance size from the security/cost tradeoff. In essence, we have a LPN variant whose security and computational cost both degrade gracefully in the subspace dimension.

### v3

@@ 5 An LPN Hardness Assumption and Performance Scaling
$\exists\epsilon<1$ such that fixing any $\delta>0$, there exists $n^{*}\in\mathbb{N},c>0$ such that for any $n>n^{*},n_{1}=\delta n$, $\mu=n^{\epsilon-1}$, no polynomial-time adversary can distinguish the distribution of tuples $(L,LH+S)$ from that of $(L,U)$ with non-negligible advantage, where $L\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n\times n_{1}}$, $H\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{1}\times\mathrm{\,poly}(n)}$, $U\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n\times\mathrm{\,poly}(n)}$, and $S\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathcal{D}_{R,\mu}^{n\times\mathrm{\,poly}(n)}$.
@@
Both $\delta$ and $\epsilon$ directly correspond to the parameters by the same names in Theorem 3.1. In particular, slightly stronger assumptions directly translate into significant efficiency improvements. For instance, if $\delta$ is allowed to be $o(1)$, we achieve sub-multiplicative overhead for $\frac{m}{n}\geq\delta$, rather than just $m=\Omega(n)$. All terms of the form $n^{\epsilon}$ correspond to $n\mu$; if we assume that $\mu(n)=n^{\epsilon-1}$ is secure for every $\epsilon>0$, we obtain the corresponding substantial efficiency improvements. [^6] Our LPN assumption is commonly believed to hold (nonuniformly) for all $\epsilon$, but our results are nontrivial for any $\epsilon<\omega-2$. Likewise, the security of our protocol depends directly on the computational hardness assumption: a super-polynomial assumption would yield a protocol with super-polynomial security.
6 A Simple Secure Delegated Matrix-Matrices Protocol
@@
We first use LPN to construct schemes for Trapdoored-Matrix and Targeted-Trapdoored-Matrix. Our approach is simple: we observe that a matrix of the form $LH+S$, where $L$ and $H$ have few columns and rows respectively and $S$ is sparse, is pseudorandom by the LPN assumption and is easy to multiply by other matrices. Moreover, given a matrix $A$, once $AL^{\top}$ has been computed, it is very efficient to compute $A(LH+S)$ as $(AL)H+AS$, even if $LH+S$ is a thin matrix (or vector).
This results in a simple protocol for Secure Delegated Matrix-Matrices, detailed in Protocol 6.
@@
A brief note on notation: in order to aid the reader in tracking which parts of expressions have already been computed, we use angled brackets around an expression, such as $\langle\mathrm{Expr}(x,y,z,\dots)\rangle$, to denote an already-computed variable whose value is equal to that of the expression $\mathrm{Expr}(x,y,z,\dots)$.
Protocol 1: Simple Client Protocols for Matrix-Matrices Inputs: Ring $R$, left matrix $A\in R^{m\times n}$, subspace dimension $n_{1}$, sparsity $\mu$. Initialization: 1. Sample LPN subspace matrix $L\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n\times n_{1}}$. 2. Sample $H\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{m\times n_{1}},S\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathcal{D}_{\mu}^{m\times n}$. 3. Compute $A^{\prime}\coloneqq HL^{\top}+S.$ 4. Compute and send $A_{\mathbf{enc}}\coloneqq A+A^{\prime}$. 5. Compute $A_{\mathbf{enc}}L,AL$. Multiplication by $B\in R^{n\times l}$: 1. Sample $G\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{1}\times l},T\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathcal{D}_{\mu}^{n\times l}$. 2. Compute $B^{\prime}\coloneqq LG+T.$ Compute and send $B_{\mathbf{enc}}\coloneqq B+B^{\prime}$. 3. Compute $AB^{\prime}=\langle AL\rangle G+AT$, $L^{\top}B_{\mathbf{enc}}$ 4. Request $A_{\mathbf{enc}}B_{\mathbf{enc}}$. 5. Compute $A^{\prime}B_{\mathbf{enc}}=H\langle L^{\top}B_{\mathbf{enc}}\rangle+SB_{\mathbf{enc}}.$ 6. Return $AB=\langle A_{\mathbf{enc}}B_{\mathbf{enc}}\rangle-\langle AB^{\prime}\rangle-\langle A^{\prime}B_{\mathbf{enc}}\rangle$
@@ 6.1 Security and Performance
The security of Protocol 6 is a direct consequence of the LPN assumption: the matrices $A^{\prime}$ and $B^{\prime}$ are constructed exactly as the matrix that the assumption holds to be pseudorandom conditioned on $L$, and the joint distribution of $(L,A+A^{\prime},B+B^{\prime})$ (from which all messages to the server can be derived) therefore leaks no information about $A$ and $B$.
Protocol 6 is moderately efficient, already allowing for client savings for large parameters. The initialization requires one $m\times n_{1}$-by-$n_{1}\times n$ matrix multiplication, one $m\times n$-by-$n\times n_{1}$ matrix multiplication, and negligible additions. An online step has similar cost; in particular, letting $a,b,c$ be the ascending sort of $m,n,l$, and $a^{\prime},b^{\prime},c^{\prime}$ be the ascending sort of $m,n_{1},l$, the cost for an online step is $O\left((a^{\prime})^{\omega-2}b^{\prime}c^{\prime}+mnl\mu\right)$, while insecure delegation has cost $O\left((a)^{\omega-2}bc\right)$. For instance, in the case $m=n,l=1$ (corresponding to matrix-vectors), we drive the client cost down from $O(n^{2})$ to $O(n(n_{1}+n\mu))$. A server’s online step merely consists of receiving two matrices of the problem sizes, computing the product, and returning it, so server computation time and round/bit efficiency of communication are optimal.
@@
For most parameter regimes (e.g. $\delta=\Theta(1),\mu=o(1)$), the dominant cost comes from the dense multiplication $HL$, and therefore varies polynomially in subspace dimension $n_{1}$. This can be greatly reduced by using LPN variants over fast matrices (e.g. sparse-LPN, Toeplitz-LPN, Quasi-cyclic-LPN [1]), but precise parameter tradeoffs for these problems are not currently well-understood, and may entail worse $\mu$. In the next subsection, we demonstrate a more advanced protocol that obtains similar or better performance (depending on the comparative sparsity assumptions of LPN and sparse-LPN) than a sparse-LPN version, but depends only on standard LPN.
7 An Improved Secure Delegated Matrix-Matrices Protocol
@@
Protocol 6 has client cost very roughly $(\mu+\delta)n$ times the cost of simply sending the inputs to the server. This is rather high. It is also not (just) an artifact of our weak parameter assumptions: the security of LPN does not degrade gracefully as its parameters are reduced. In particular, once $\delta\mu\leq\frac{1}{n}$, straightforward attacks from subsampling and Gaussian elimination begin to become efficient. This imposes a minimum multiplicative overhead of $\Theta(\sqrt{n})$ if super-polynomial security is desired. In this section, we present a more complicated protocol that bypasses this barrier to improve performance.
@@
The client cost of Protocol 6 is dominated by the costs of the multiplications $AL$ and $(AL)G$. There is a natural solution: recursively call the server to compute these, while being careful to continue hiding $A$ and $G$. Additionally, computations where both matrices are public (for instance, those of the form $L(B+B^{\prime})$) may be offloaded to the server. These changes result in a fairly complicated recursive protocol with several disadvantages, including a number of rounds growing exponentially in the recursion depth and substantial duplication of work, but this protocol in turn may be greatly optimized and simplified. We detail the resulting protocol for secure delegated matrix-matrices in Protocol 7 (and a non-interactive verification subroutine based on Freivalds’ algorithm [19] in Protocol 7), and spend the rest of this section on exposition and showing that the protocol fulfills Theorem 3.1.
Protocol 2: Checking Partial Products Inputs: Ring $R$, matrices $M_{i}\in R^{a_{i}\times a_{i+1}}\forall i\in[d]$, matrices $P_{i}\in R^{a_{1}\times a_{i+1}}\forall i\in\{2,3,\dots,d\}$, number of product checks $\lambda^{\prime}$. 1. Let $P_{1}=M_{1}$. 2. For each $i\in\{2,\dots,d\}$ 2.1. Sample $X\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{a_{i+1}\times\lambda^{\prime}}$. $2.$ If $P_{i}X\neq P_{i-1}(M_{i}X)$, return False. 3. Return True.
Protocol 3: Improved Client Protocols for Matrix-Matrices Inputs: Ring $R$, left matrix $A\in R^{m\times n}$, subspace dimensions $(n=n_{0}),n_{1},\dots,n_{d}$, sparsities $\mu_{1},\dots,\mu_{d}$, number of product checks $\lambda^{\prime}$. Initialization: 1. Sample and send LPN subspace matrices $L_{i}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{i-1}\times n_{i}}\forall i\in[d]$. 2. For each pair $(i,j)\in\{0,1,\dots,d\}^{2}$, request the partial product $L_{j}^{\top}L_{j-1}^{\top}\dots L_{1}^{\top}L_{1}L_{2}\dots L_{i}$. Check these via Protocol 7, and abort if any are incorrect. 3. Sample $H\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{m\times n_{d}},S_{i}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathcal{D}_{\mu_{i}}^{m\times n_{i-1}}\forall i\in[d]$. 4. Compute $A^{\prime}\coloneqq H\langle L_{1}L_{2}\dots L_{d}\rangle^{\top}+\sum_{i=0}^{d-1}S_{i+1}\langle L_{1}L_{2}\dots L_{i}\rangle^{\top}.$ 5. Compute and send $A_{\mathbf{enc}}\coloneqq A+A^{\prime}$. 6. Request and check via Protocol 7 the partial products $A_{\mathbf{enc}}L_{1}$, $A_{\mathbf{enc}}L_{1}L_{2}$, …, $A_{\mathbf{enc}}L_{1}L_{2}\dots L_{d}$. Abort if any are incorrect. 7. Compute, for every $i\in[d]$, $A^{\prime}L_{1}L_{2}\dots L_{i}$ $=H\langle L_{d}^{\top}L_{d-1}^{\top}\dots L_{1}^{\top}L_{1}L_{2}\dots L_{i}\rangle$ $\quad+\sum_{j=0}^{d-1}S_{j+1}\langle L_{j}^{\top}L_{j-1}^{\top}\dots L_{1}^{\top}L_{1}L_{2}\dots L_{i}\rangle$ $AL_{1}L_{2}\dots L_{i}$ $=\langle A_{\mathbf{enc}}L_{1}L_{2}\dots L_{i}\rangle-\langle A^{\prime}L_{1}L_{2}\dots L_{i}\rangle.$ Multiplication by $B\in R^{n\times l}$: 1. Sample $G\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{d}\times l},T_{i}\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathcal{D}_{\mu_{i}}^{n_{i-1}\times l}\forall i\in[d]$. 2. Compute $B^{\prime}\coloneqq\langle L_{1}L_{2}\dots L_{d}\rangle G+\sum_{i=0}^{d-1}\langle L_{1}L_{2}\dots L_{i}\rangle T_{i+1}.$ Compute and send $B_{\mathbf{enc}}\coloneqq B+B^{\prime}$. 3. Compute $AB^{\prime}=\langle AL_{1}L_{2}\dots L_{d}\rangle G+\sum_{i=0}^{d-1}\langle AL_{1}L_{2}\dots L_{i}\rangle T_{i+1}$. 4. Request $A_{\mathbf{enc}}B_{\mathbf{enc}}$ and partial products $(L_{1})^{\top}B_{\mathbf{enc}}$, $(L_{1}L_{2})^{\top}B_{\mathbf{enc}}$, …, $(L_{1}L_{2}\dots L_{d})^{\top}B_{\mathbf{enc}}$. 5. Compute $A^{\prime}B_{\mathbf{enc}}=H\langle(L_{1}L_{2}\dots L_{d})^{\top}B_{\mathbf{enc}}\rangle+\sum_{i=0}^{d-1}S_{i+1}\langle(L_{1}L_{2}\dots L_{i})^{\top}B_{\mathbf{enc}}\rangle.$ 6. Return $AB=\langle A_{\mathbf{enc}}B_{\mathbf{enc}}\rangle-\langle AB^{\prime}\rangle-\langle A^{\prime}B_{\mathbf{enc}}\rangle$
7.1 Protocol Structure
The structure of Protocol 7 closely parallels that of Protocol 6. The key difference is that pseudorandom matrices are now constructed via iteration of the LPN construction; for instance,
$B^{\prime}$ $\coloneqq L_{1}L_{2}\dots L_{d}G+\sum_{i=0}^{d-1}L_{1}L_{2}\dots L_{i}T_{i+1}$
$=(T_{1}+L_{1}(T_{2}+L_{2}(T_{3}+L_{3}(\dots(T_{d}+L_{d}G))))).$
This change makes the client-side dense multiplications substantially smaller, improving efficiency. It does, however, require precomputation of various products of the LPN subspace matrices and $A$. Luckily, these products can themselves be offloaded to the server (including by a partially recursive masking trick, as seen on line 7), and owing to the thinness of the subspace matrices, have low cost relative to the target multiplications $AB_{i}$. The client must check the honesty of the server on these multiplications; Protocol 7, a slight generalization of Freivalds’ algorithm [19], does so efficiently.